OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -4 + 5 + 14 + 23 + ... + 131

OpenStudy (anonymous):

First element is -4 Second elemnt is -4 + 9 Third element is -4 + 9 + 9 Nth element is -4 + 9*(n-1) Lets look at 131: 131 = -4 + 9*(n-1) --> n-1 = 15 --> n=16 So 131 is the 16th element \[\sum_{1}^{16} -4+9(n-1) = \sum_{1}^{16} -4+9n-9 = \sum_{1}^{16} 9n-13\]

OpenStudy (anonymous):

Okay, that is the same answer I obtained, however it is not one of my choices.

OpenStudy (anonymous):

Send your choices, let me have look at them

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):



OpenStudy (anonymous):

Wow, I accidentally closed this.

OpenStudy (anonymous):

\[\sum_{n=1}^{16} 9n-13 \rightarrow \sum_{k=0}^{15} 9k-4 \rightarrow \sum_{n=0}^{15} 9n-4\] There is a secret n in the top of the summation symbol. So 16 becomes 15

OpenStudy (anonymous):

Okay, Thank You so much !